OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1

OpenStudy (jdoe0001):

 \(\bf tan(\theta)=\cfrac{y}{x}\qquad tan\left(-\frac{3\pi}{4}\right)=\cfrac{y}{-1}\implies -1\cdot tan\left(-\frac{3\pi}{4}\right)=y\)